OpenStudy (anonymous):

Consider the level surface given by 2x2+5y2+z2=4. Match the slices with their correct plots below.

OpenStudy (tkhunny):

That's a lovely ellipsoid, so they better all look like ellipses.

OpenStudy (anonymous):

So how to match

OpenStudy (tkhunny):

Well, how's your Analytic Geometry? Start with this: \(\dfrac{x^{2}}{2} + \dfrac{y^{2}}{4/5} + \dfrac{z^{2}}{4} = 1\) Set x = 0 and get: \(\dfrac{y^{2}}{4/5} + \dfrac{z^{2}}{4} = 1\) This, you should recognize as an ellipse with the axis in the y-direction of \(2/\sqrt{5}\) and the axis in the z-direction of \(2\). Do you see one of those? Set y = 0 and get: \(\dfrac{x^{2}}{2} + \dfrac{z^{2}}{4} = 1\) This, you should recognize as an ellipse with the axis in the x-direction of \(\sqrt{2}\) and the axis in the z-direction of \(2\). Do you see one of those? Continue with this sort of investigation.

OpenStudy (anonymous):

It's look like football :DD

OpenStudy (tkhunny):

Yes, yes it does. But today, we're projecting the football onto a plane and examining the cross sections.

OpenStudy (anonymous):

Thank you so much, tkhunny! I got the answer!